OpenStudy (cloverloveee16):

Its been a a year or 2 since I have takin math. How do you do equations (y=3x-3). I can't really explain, but just explain anything that has this.

OpenStudy (phi):

y=3x-3 is the equation of a line people learn to plot lines on graph paper how to find y if they know x. for example, if x is 0, then y = 3*0-3 or just y= -3 people learn that the "slope" is 3 and the "y-intercept" is -3

OpenStudy (cloverloveee16):

Ok so can you please explain f(x) things.?

OpenStudy (phi):

when people write f(x) that is short for: there is a rule (we name it "f"), that tells us how to calculate an answer that depends on "x" x is a "place-holder" in the rule, and is replaced by a number when we do the calculation

OpenStudy (phi):

for example, say we have f(x) = x+1 the rule (named "f") says: given a number add 1 to it f(1) is how we say what number we want to start with: f(1) = 1+1 (we replace the x with 1) f(1) = 2 that means when we use rule f with x=1 we get 2 as our answer.

OpenStudy (cloverloveee16):

Thank you so much @phi. You are my hero ! :) <3

OpenStudy (phi):

people usually call f a "function" instead of a rule but it means pretty much the same thing. But to keep things simple, a function (or rule) is only allowed to give one number for each x Here is a rule that is not a function: sqrt(x) that is not a function because sqr(4) = 2 or sqr(4) = -2 (both -2 and 2 work) but we are not allowed to have two different answers. so f(x) = sqr(x) is not allowed (unless we say: only the positive values are allowed... then we can use it as a function)